OpenStudy (anonymous):

A manufacture has been selling 1400 television sets a week at 420 each. A market survey indicates that for each 21 rebate offered to a buyer, the number of sets sold will increase by 210 per week. a) Find the demand function p(x), where x is the number of the television sets sold per week. p(x)= b) How large rebate should the company offer to a buyer, in order to maximize its revenue? c) If the weekly cost function is 98000+140x, how should it set the size of the rebate to maximize its profit?

OpenStudy (anonymous):

Revenue: R(x) = P(x) * x = ( 420 - 21x ) ( 1,400 + 210 x) Maximize Revenue: R'(x) = 0 => x ≈ 6.67 => Rebate size = 6.67 * $21 = $ 140 Profit : P(x) = R(x) - C(x) = ( 420 - 21x ) ( 1,400 + 210 x) - ( 98000+140x) .... P'(x) = 0 -> x = ... => Rebate size: ..